Give the dimension of the water tank, use its diagonal for the height.l = 125 cm and width = 129 cm. How to determine the diagonal of the rectangular water tank?
Answer by Theo(13342)

the length is 125 cm.
the width is 129 cm.
the diagonal and the length and the width form a right triangle with one leg = 125 cm and the other leg = 139 cm.
the diagonal is equal to sqrt(125^2 + 129^2) = 179.6273921 cm

you are using the pythagorian formula of a right triangle which is:
H^2 = L1^2 + L2^2
H is the hypotenuse
L1 is one of the legs of the right triangle.
L2 is the other leg of the right triangle.
take the square root of H^2 and you have H.